OpenStudy (anonymous):

sqrt(x - 1) = x - 7? please help!

hero (hero):

Hint: Square both sides to get x - 1 = (x - 7)^2

OpenStudy (saifoo.khan):

Take square on both sides, \[(x-1) = (x-7)^2\]

OpenStudy (anonymous):

especially with these, check your answer(s)...:)